Question 2: solve the differential equations a) (xy - y)dx + - x)dy = 0
Answer #1

I m solving this in two different ways.Rate it.@ Giren (my-y) dx + ( 2 - kdy Comparing 0 with Mdatindy =0 , we get - Many y; N= = 2 x is an ex-1 l n = x=1 since 2M = SN , to is exact Jmdr = f (my-y) dx Ty in supposed = mong - xy to be and Indy =) ( 22 - 2) dy = aryl - ny complete primitive in giv